OpenStudy (anonymous):

If f′ is continuous, f(6)=0, and f′(6)=13, evaluate

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} (f(6+3x)+f(6+5x))/x\]

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

Gah, haven't done this in forever, and I gotta run - sorry :/ @mashy @shamil98 Can you help?

OpenStudy (anonymous):

no worries

OpenStudy (anonymous):

@ranga @zepdrix

OpenStudy (anonymous):

\[f \prime \left( c \right)=\lim_{x \rightarrow c}\frac{ f \left( x \right)-f \left( c \right) }{ x-c }\]

OpenStudy (anonymous):

\[If f \left( x \right) is continuous at any point a then \lim_{x \rightarrow a}f \left( x \right)=f \left( a \right)\]

OpenStudy (anonymous):

so how should i answer the question

OpenStudy (ranga):

@muzzammil.raza Have you been taught L'Hopital's Rule yet?

OpenStudy (anonymous):

i don't think he is supposed to use L'hospital here :P..

OpenStudy (ranga):

\[ \lim_{x \rightarrow 0} \frac{f(6+3x)+f(6+5x) }{ x } \text{ = } \lim_{x \rightarrow 0} \frac{ f(6+3x)}{ x } \text{ + }\lim_{x \rightarrow 0} \frac{ f(6+5x)}{ x } \]I will do the first limit here: lim x->0 f(6+3x) / x Write f(6+3x) / x as { f(6+3x) - f(6) } / x {since f(6) = 0} { f(6+3x) - f(6) } / x = 3{ f(6+3x) - f(6) } / 3x Let u = 3x. As x->0, u->0 The limit becomes: lim u->0 3 { f(6 + u) - f(6) } / u = 3 * lim u->0 { f(6 + u) - f(6) } / u = 3 * f'(6) = 3 * 13 = 39